Assume that there are 538, 584 males 50 and older alive with CRC, and 31,346 males younger than 50 alive with CRC. What are the prevalence rates (per 100,00) for each age group?
Answer #1

CRC stands for colorectal cancer. Basically it is more common in individual more than 50 years of age.

Prevelance rate for male 5 yrs old or above -- number of lives / 10000. =538584/10000= 53.8

For male younger than 50 year = 31346/10000=3.13

Hence chances for CRC are more in made above 50 years of age.